OpenStudy (anonymous):

Which of the following would be part of the correct solution for completing the square of the equation X^2+10x=3? A) (x+5)^2=13 B) (x+5)^2=28 C) (X-5)^2= -22 D) ( x+25)^2=28

OpenStudy (anonymous):

Sorry about that earlier reply. The answer is not D. The perfect square polynomial starting x^2+10x is x^2+10x+25, with it factoring to (x+5)^2. Therefore, add 25 to each side of the equation to get the left side into the form that you want. Then you have x^2+10x+25=3+25. Factoring and simplifying gives you (x+5)^2=28. The answer is B

OpenStudy (anonymous):

how did you get 25 in the problem

OpenStudy (anonymous):

can you please draw it please

OpenStudy (anonymous):

SORRY for the trouble

OpenStudy (anonymous):

I tried and did what you said but I end up with X^2+10x=26

OpenStudy (anonymous):

i mean 28 oops

OpenStudy (anonymous):

 I don't really know what else to draw. If the coefficient on your x^2 term is 1, then the coefficient on your x term is going to be 2*(whatever number is in your squared polynomial). For example, if you were given: "x^2 + 6x" is the start of a perfect square polynomial. Finish it. Then you would just take 6/2 = 3. The perfect square factored is (x+3)^2. So the perfect square would be x^2+6x+9